OpenStudy (anonymous):

A runner starts from rest and has an acceleration of 3 m/s2. How fast is she running after 1.1 seconds?

OpenStudy (anonymous):

you can use: \[v(t)=v_0+at\] - t and a are given - v0 should be the initial speed, what calue do you think it takes for this problem?

OpenStudy (anonymous):

v = u + at v = 0 + 3x1.1 v=3.3ms-1